In Mexico, avocados are 50 pesos per kilogram. What is the cost in cents of an avocado that weighs 0.40 lb if the exchange rate is 10 pesos to the dollar?
Answer by richwmiller(17219)

one kilo is 2.2 lbs
.4 lbs divided by 2.2 =
2/5 divided by 2.2 (11/5)
2/5*5/11=2/11
50/10=5 dollars per kilo
2/11*5=10/11 dollars =10/11*100 about 91 cents